OpenStudy (anonymous):

Find the exact value of cos(arcsin(one fourth)). For full credit, explain your reasoning.

OpenStudy (psymon):

So what this question basically asks is find the cosine of a triangle that has a sin value of 1/4. So we want a triangle that has a sin (opposite/hypotenuse) of 1/4 and then find thecosine of that triangle. So doing that we can draw this:  Now using that, find the cosine :3

OpenStudy (anonymous):

sin^-1(1/4)=x right?

OpenStudy (psymon):

No, we find x through pythagorean theorem.

OpenStudy (anonymous):

ok \[1^{2} + x ^{2} = 4^{2}\] \[x^{2} = 16-1\] \[x ^{2}=15\] \[x=\sqrt{15}\]

OpenStudy (psymon):

Correct. So now with all 3 sides, what is the cosine of that triangle?

OpenStudy (anonymous):

cos= \[\frac{ \sqrt{15} }{ 4}\]

OpenStudy (psymon):

That's your answer :3

OpenStudy (anonymous):

Thank you very much :)

OpenStudy (psymon):

Yep, np :3